# Contextual Status Updating: How to Maximize Situational Awareness?

### v1

@@ Appendix E Special Case: Single-Source, Single-Channel
Problem (57) is an MDP that can be solved by Dynamic programming [17]. The optimal policy associated with agent $1$ satisfies the following Bellman optimality equation:
$J_{1}(\delta,x)=$ $H_{L}(Y_{1,\delta}|X_{1,0}=x)-\mathsf{L}_{1,\text{opt}}$
$+\min\{J_{1}(\delta+1,x),(1-p_{1})J_{1}(\delta+1,x)$
$+p_{1}\mathbb{E}[J_{1}(1,X_{1,0})|X_{1,-\delta}=x]\},$ (58)
@@
$J_{1,k+1}(\delta,x)=$ $H_{L}(Y_{1,\delta}|X_{n,0}=x)-\mathsf{L}_{n,\text{opt}}+$
$\min\{J_{1,k}(\delta+1,x),(1-p_{1})J_{1,k}(\delta+1,x)$
$+p_{1}\mathbb{E}[J_{1,k}(1,X_{1,0})|X_{1,-\delta}=x]\},$ (59)
which converges to $\lim_{k\to\infty}J_{1,k}=J_{1}$ for any $J_{1,0}$. After some rearrangements, we can write (E) as
$J_{1,k+1}(\delta,x)=$ $H_{L}(Y_{1,\delta}|X_{1,0}=x)-\mathsf{L}_{1,\text{opt}}+$
$J_{1,k}(\delta+1,x)+p_{1}\min\{0,-J_{1,k}(\delta+1,x)$
$+\mathbb{E}[J_{1,k}(1,X_{1,0})|X_{1,-\delta}=x]\}.$ (60)
@@
See Appendix F. ∎
Using Lemma 4, we get that for any $k$, $J_{1,k}(\delta+1,x)\geq\mathbb{E}[J_{1,k}(1,X_{1,0})|X_{1,-\delta}=x]$ which implies that the penalty for not sending at iteration step $k$ is higher than sending. Therefore, taking the active action is beneficial to reduce the penalty. One interesting observation from (E) is that each time a packet is successfully delivered with probability $p_{1}$, a new piece of information about the sensor signal value is added with the existing information $(X_{1,t-\delta}=x)$ (see the term $\mathbb{E}[J_{1,k}(1,X_{1,0})|X_{1,t-\delta}=x]$ in (E)). This new information plays a crucial role in reducing the system penalty and hence benefits the system through sending. This completes the proof of Theorem 3.
Appendix F Proof of Lemma 4
@@
$\geq$ $\!\sum_{x\in\mathcal{X}}\!P(X\!=\!x|Z\!=\!z)\min_{a\in\mathcal{A}}\!\sum_{y\in\mathcal{Y}}\!P(Y\!=\!y|X\!=\!x,Z\!=\!z)L(y,a),$ (63)
@@
$H_{L}(Y|Z=z)$
$\geq$ $\sum_{x\in\mathcal{X}}P(X=x|Z=z)\min_{a\in\mathcal{A}}\mathbb{E}[L(Y,a)|X=x,Z=z].$ (64)

### v2

@@ Appendix F Special Case: Single-Source, Single-Channel
Problem (60) is an MDP that can be solved by Dynamic programming [17]. The optimal policy associated with agent 111 satisfies the following Bellman optimality equation:
$J_{1}(\delta,x)=H_{L}(Y_{1,\delta}|X_{1,0}=x)-\mathsf{L}_{1,\text{opt}}+\min\{J_{1}(\delta+1,x),(1-p_{1})J_{1}(\delta+1,x)+p_{1}\mathbb{E}[J_{1}(1,X_{1,0})|X_{1,-\delta}=x]\},$ (61)
@@
$J_{1,k+1}(\delta,x)=H_{L}(Y_{1,\delta}|X_{n,0}=x)-\mathsf{L}_{n,\text{opt}}+\min\{J_{1,k}(\delta+1,x),(1-p_{1})J_{1,k}(\delta+1,x)+p_{1}\mathbb{E}[J_{1,k}(1,X_{1,0})|X_{1,-\delta}=x]\},$ (62)
which converges to $\lim_{k\to\infty}J_{1,k}=J_{1}$ for any $J_{1,0}$. After some rearrangements, we can write (62) as
$J_{1,k+1}(\delta,x)=H_{L}(Y_{1,\delta}|X_{1,0}=x)-\mathsf{L}_{1,\text{opt}}+J_{1,k}(\delta+1,x)+p_{1}\min\{0,-J_{1,k}(\delta+1,x)+\mathbb{E}[J_{1,k}(1,X_{1,0})|X_{1,-\delta}=x]\}.$ (63)
@@
See Appendix G. ∎
Using Lemma 4, we get that for any $k$, $J_{1,k}(\delta+1,x)\geq\mathbb{E}[J_{1,k}(1,X_{1,0})|X_{1,-\delta}=x]$ which implies that the penalty for not sending at iteration step $k$ is higher than sending. Therefore, taking the active action is beneficial to reduce the penalty. One interesting observation from (63) is that each time a packet is successfully delivered with probability $p_{1}$, a new piece of information about the sensor signal value is added with the existing information $(X_{1,t-\delta}=x)$ (see the term $\mathbb{E}[J_{1,k}(1,X_{1,0})|X_{1,t-\delta}=x]$ in (62)). This new information plays a crucial role in reducing the system penalty and hence benefits the system through sending. This completes the proof of Theorem 3.
Appendix G Proof of Lemma 4
@@
$\geq\!\sum_{x\in\mathcal{X}}\!P(X\!=\!x|Z\!=\!z)\min_{a\in\mathcal{A}}\!\sum_{y\in\mathcal{Y}}\!P(Y\!=\!y|X\!=\!x,Z\!=\!z)L(y,a)$ (67)
@@
$H_{L}(Y|Z=z)\geq\sum_{x\in\mathcal{X}}P(X=x|Z=z)\min_{a\in\mathcal{A}}\mathbb{E}[L(Y,a)|X=x,Z=z].$ (68)